OpenStudy (anonymous):

Hey guys , about proving integral secx = ln secx+tanx since ln is an abs value , my prof said that he don't need how to get the answer , he just wants me to take the derivative of secx+tanx , but he also said I should consider it when it's positive and negative ? how can I solve it if it's negative ?

OpenStudy (amistre64):

ln(-a) tends to be a complex value, if memory serves

OpenStudy (anonymous):

The problem is it's been about 4 years since I've taking calculus that involves complex numbers. :(

OpenStudy (anonymous):

i'm not sure about positive and negative , but when I did it I took the derivative of secx+tanx over secx+tanx which will be secx , i'm not sure if this is enough or what :D

OpenStudy (amistre64):

\[\int sec(x)\frac{sec(x)+tan(x)}{sec(x)+tan(x)}dx\] \[\int \frac{sec^2(x)+sec(x)tan(x)}{tan(x)+sec(x)}dx\] the top is the derivative of the bottom .... which ideally ints up to ln|u| sec+tan has 0 and negative parts .... so i am not sure what the added +- is spose to suggest

OpenStudy (anonymous):

well , I'm confused :D I remember him saying since it's abs check both negative and positive I already simplified after taking the derivative of ln abs u and got secx .

OpenStudy (amistre64):

it seems that log(-x) = log(x) + i pi

OpenStudy (anonymous):

and i represent an imaginary ? but how could I SOLVE IT ?

OpenStudy (amistre64):

i pi is just a constant, but then i dont know if this is the right track to take :/

OpenStudy (anonymous):

well , I'm kinda confused right now , first I thought he was just trying to explain how antiderivatives work , but when I remembered saying look for the negative and positive I got confused .

OpenStudy (amistre64):

ask him to explain himself better :)

OpenStudy (anonymous):

well I'm pretty sure he'll just say the same things , look for the negative and positive side. also I remember when we got an abs we use what's called , sorry I don't remember it's name in english , but it's " -b+or- sqrt(b^2-4ac) and check for the zeros or something to see if it's positive or negative

OpenStudy (anonymous):

OK GUYS , let me ask this , if I done integrating for example my result was ..: \[\ln \left| something \right|\] how can we determine if it was negative to keep the abs , or positive to remove it and replace it by (something) ?

OpenStudy (joannablackwelder):

My experience is that you keep the absolute value if what is inside could be negative. You can drop it if it could never be negative.

OpenStudy (joannablackwelder):

Such as you could drop the absolute value for ln(x^2) since it can never be negative.

OpenStudy (anonymous):

well , in some cases it could be negative , but I think there's a better way to solve and see if it's negative or positive.

OpenStudy (joannablackwelder):

Maybe this link will be helpful http://www.math.ubc.ca/~feldman/m121/secx.pdf

OpenStudy (joannablackwelder):

Since it is absolute value, when the inside is negative, take the absolute value before taking the ln.

OpenStudy (anonymous):

Thanks buddy , I'll check with my prof and see if he could explain. Thanks all of u.